# Addressing Hallucinations with RAG and NMISS in Italian Healthcare LLM Chatbots

### v2

@@ 3.1 RAG Architecture
RAG models combine the strengths of retrieval-based and generative language models to address issues like hallucinations and factual inaccuracies by integrating a retrieval mechanism to ground responses in relevant information [24].
I formally define the problem as follows. Let $q$ represent the query, $c$ the context information retrieved from external sources, and $r$ the output sequence generated by the model. The architecture consists of two components: (i) a retriever $p_{\eta}(c|q)$ with parameters $\eta$, which selects a top-K set of documents $k$ based on the query $q$, and (ii) a generator $p_{\theta}(r_{i}|q,c,r_{1:i-1})$, with parameters $\theta$, which generates each token $r_{i}$ conditioned on the query $q$, retrieved context $c$, and previously generated tokens $r_{1:i-1}$.
@@
This RAG-sequence approach ensures that the response remains consistent by using the same retrieved document to generate each token, maintaining coherence throughout the sequence. Here, $N$ represents the length of the output sequence. This architecture is particularly advantageous for domains like healthcare, where factual accuracy is critical. By leveraging retrieval from medical documents, guidelines, or research articles, the RAG model ensures that responses are accurate, minimizing the risks associated with misinformation [10]. In this paper, I employ the retrieval and generation models separately and combine them for inference; the workflow is detailed in Section 4.2.
@@ 3.2 Evaluation Metrics
The metrics described in this section are used to evaluate the quality of responses generated by the system, referred to as the System Summary (SS), against the ground truth responses, referred to as the Reference Summary (RS), which are human-generated responses in this study. Both SS and RS are treated as sets of n-grams for all metrics. Formally, let $S_{\text{SS}}=\{s_{1},s_{2},\dots,s_{l}\}$, where $s_{i}$ is an n-gram in the SS, and $S_{\text{RS}}=\{r_{1},r_{2},\dots,r_{m}\}$, where $r_{j}$ is an n-gram in the RS.
ROUGE [26] evaluates responses based on precision, recall, and F1-score. The text is pre-processed before calculating these scores [43]. More formally, ROUGE-N measures the overlap of n-grams between the SS and RS. The precision, recall, and F1-score for ROUGE-N are defined as follows:
@@
[…]
where $S_{\text{SS}}$ is the set of n-grams in the SS, $S_{\text{RS}}$ is the set of n-grams in the RS, $|S_{\text{SS}}\cap S_{\text{RS}}|$ is the cardinality of the intersection of $S_{\text{SS}}$ and $S_{\text{RS}}$, $|S_{\text{SS}}|$ is the cardinality of the set $S_{\text{SS}}$, and $|S_{\text{RS}}|$ is the cardinality of the set $S_{\text{RS}}$. Precision measures the accuracy of the generated text, while recall evaluates the coverage of the information. Low precision suggests hallucination, while low recall indicates a coverage issue.
Another variant used here is ROUGE-L, which measures the Longest Common Subsequence (LCS) between the SS and RS. The LCS represents the longest subsequence in both SS and RS while preserving word order, though the words do not need to appear consecutively [7].
BLEU [36] is designed to evaluate the quality of SS responses by comparing them against one or more RS answers. BLEU accounts for variability in word choice and word order found in different, yet valid, translations of the same source sentence. It quantifies the overlap by measuring how many n-grams from the SS are present in the RS. The core of the BLEU metric is the modified n-gram precision, which addresses the over-generation problem by treating each reference word as exhausted once a matching candidate word is found.
@@
The final BLEU score is computed by taking the geometric mean of the modified n-gram precisions (up to length $N$, typically set to 4), and multiplying by the brevity penalty:
@@
METEOR [4] addresses several limitations of BLEU by emphasizing word-to-word alignment between the SS and RS, accounting for exact matches, stems, and synonyms. METEOR computes the score by comparing the SS to each RS and reporting the highest score. Additionally, METEOR introduces a Discontiguous Matches Penalty (DP), based on the number of matched chunks.
Let $T$ represent the number of contiguous chunks of matched unigrams between SS and RS. For a set of matched unigrams $V=\{v_{1},v_{2},\dots,v_{k}\}$, a chunk is defined as a maximal sequence of consecutive matches $v_{i},v_{i+1},\dots,v_{i+t}$, such that $v_{i+t}=v_{i}+t$ for all $t\geq 0$.
The function $T(V,S_{\text{SS}})$, which counts the number of contiguous chunks, is then formalized as:
$T(V,S_{\text{SS}})=\left|\{\{v_{i},v_{i+1},\dots,v_{i+t}\}\subseteq V\mid v_{i+t}=v_{i}+t\}\right|$ (8)
This function returns the number of contiguous chunks of matched unigrams in $S_{\text{SS}}$. The Discontiguous Matches Penalty (DP) is then computed as:
$\text{DP}=0.5\cdot\left(\frac{T(V,S_{\text{SS}})}{|V|}\right)$ (9)
where $|V|$ is the number of matched unigrams. The penalty increases as the number of chunks grows, penalizing fragmented matches. The final METEOR score is adjusted by applying DP to the $F_{\text{mean}}$ score:
$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{DP})$ (10)
BERTSCORE is an evaluation metric for text generation that leverages pre-trained BERT embeddings [58]. Unlike traditional metrics that rely on exact n-gram matches, BERTSCORE measures the semantic similarity between tokens in the System Summary (SS) and Reference Summary (RS) by comparing their contextual embeddings. This method captures the semantic meaning even when the exact words differ. BERTSCORE generates contextual embeddings for the tokens in both the SS and RS, where each embedding reflects the surrounding context of the word. The similarity between tokens is computed using cosine similarity. For each token in the RS, BERTSCORE identifies the most similar token in the SS, and vice versa, to compute precision and recall.
Given a RS sentence $x=\{x_{1},\ldots,x_{k}\}$ and a SS sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$, the token embeddings are represented as $\{e(x_{1}),\ldots,e(x_{k})\}$ and $\{e(\hat{x}_{1}),\ldots,e(\hat{x}_{l})\}$, where $e(\cdot)$ represents the contextual embedding for each token. The similarity between two embeddings is computed using cosine similarity. Precision ($P_{\text{BERT}}$) and recall ($R_{\text{BERT}}$) are calculated as follows:
@@
3.3 NMISS
@@ 3.3.1 Rationale
The metrics discussed above evaluate the quality of text summaries and machine-generated responses by comparing them to reference answers. While effective, this method overlooks the possibility that non-matching words or n-grams might still be valuable if they align with the question’s broader context. This observation prompts an investigation into whether the words or phrases that do not match the reference could still be relevant compared to the broader context in which the response is generated. LLMs often produce responses beyond the strict instructions, frequently adding extra content, as shown in Table 1. This additional content can manifest in two ways: as hallucinations, where the model generates information that is factually inaccurate or irrelevant, or as meaningful elaborations, where the model extends the response in a contextually relevant manner.
@@
By accounting for these non-matching elements, I can more effectively evaluate the performance of LLMs, distinguishing between irrelevant or hallucinated information and valuable, contextually appropriate elaborations. This approach offers a more nuanced understanding of the model’s output.
The NMISS scoring system was developed to capture the potential value in these unmatched words by evaluating whether they align with the contextual information provided. This rationale stems from the observation that LLMs often do not strictly adhere to the given instructions, incorporating additional content into their answers. This content may either be a hallucination or a contextually relevant elaboration that enriches the response.
@@ 3.3.2 Formulation
Let =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ represent the set of reference answers, $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$ the set of system-generated answers, and $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ the set of context texts associated with each reference answer. Define the set of matching tokens between system-generated answers and reference answers as $\mathcal{P}(s_{i},r_{i})$, and let $\mathcal{N}(s_{i},r_{i})=s_{i}\setminus\mathcal{P}(s_{i},r_{i})$ represent the tokens in $s_{i}$ that do not match $r_{i}$. Finally, let $\mathcal{A}(s_{i},c_{i})=\mathcal{N}(s_{i},r_{i})\cap c_{i}$, which represents the set of unmatched tokens in $s_{i}$ that match the context $c_{i}$. I aim to define a function
@@
that evaluates how well system-generated answers align with both reference answers and the context. First, I define a reference-based evaluation function $f_{\text{ref},i}(s_{i},r_{i})=\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$, and a context-based evaluation function $f_{\text{cxt},i}(s_{i},c_{i})=\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$, which computes the score based on tokens matching the context. The NMISS evaluation function is defined as:
@@
To safeguard the integrity of standard scoring metrics, NMISS implements two key mechanisms: a weighted scheme and a maximum function. These mechanisms are essential for balancing reference-based and context-based contributions, addressing challenges such as score underestimation or overestimation when processing contextual information. The weighted scheme ensures a fair combination of reference-based and context-based scoring, preventing a disproportionate reliance on one over the other. Specifically, $\lambda_{1}$ and $\lambda_{2}$ control the relative influence of reference overlap and contextual alignment. This mechanism helps NMISS capture instances where a system-generated response extends beyond the reference but remains faithful to the context, distinguishing it from hallucinated content. The maximum function further enhances robustness by preventing NMISS from being unfairly penalized when context length inflates the denominator. In recall-sensitive metrics, an inflated denominator—such as the total number of n-grams in the context—could lower the score even if the additional content is meaningful. The formulation in Equation 16 ensures that context-based contributions only positively influence the score when they genuinely add value. In other words, if context matching introduces more noise than useful content, NMISS defaults to the reference-based score. This design prevents cases where large but irrelevant contexts artificially suppress the metric. This approach is particularly advantageous in tasks like question-answering, where the model’s response might not strictly match the reference but still be valid given the context. By integrating these mechanisms, NMISS ensures coherence in scoring, making it a more flexible and context-aware evaluation metric. The detailed implementation is outlined in Algorithm 1.
@@
1:=$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$, the set of reference answers
2:$\mathcal{S}=\{s_{1},s_{2},\ldots,s_{n}\}$, the set of system-generated answers
3:$\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$, the set of context texts
4:=$\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$, the set of NMISS scores for a metric $\mathcal{X}$
@@
8:     $\mathcal{P}(s_{i},r_{i})\leftarrow s_{i}\cap r_{i}$triangleright Matched tokens
9:     $\mathcal{N}(s_{i},r_{i})\leftarrow s_{i}\setminus\mathcal{P}(s_{i},r_{i})$triangleright Unmatched tokens
@@
13:     $\mathcal{A}(s_{i},c_{i})\leftarrow\mathcal{N}(s_{i},r_{i})\cap c_{i}$triangleright Context-matched tokens
@@ 4.1 Data
Data comprises text-based news in Italian, collected from various sources, and all focused on the health domain within Italy. This collection encompasses regional updates, significant scientific developments, national headlines, and news on research and development, providing comprehensive coverage of health-related issues across the country. I collected 126,470 news articles from 2010 to 2024; their frequency is shown in Figure 1. A pre-processing step was employed to clean the raw text of HTML tags and other unwanted artifacts to ensure data quality. Token statistics are calculated to understand the overall structure of the data and are shown in Table 2. On average, each article contains approximately 348 tokens, with a standard deviation of 264 tokens, indicating some variability in article length. The shortest article consists of just 6 tokens, while the longest contains up to 5,594 tokens. The median article length is 284 tokens, with the interquartile range spanning from 195 to 421 tokens, reflecting a diverse range of article sizes.
@@
Next, I developed a list of 100 questions related to public health. Questions range from very generic to specific or ambiguous, leading to different levels of interpretation ease. The questions are categorized into three levels of understanding, based on the objective criterion of question complexity, measured by the average question length as a proxy for how detailed the question is (low: 30 characters, mid: 40 characters, high: 50 characters). Additionally, all questions were reviewed by a domain expert to ensure accuracy and relevance. All questions have been answered manually; Table 3 shows an example of questions.

### v3

@@ 3.1 RAG Architecture
RAG models combine the strengths of retrieval-based and generative language models to address issues such as hallucinations and factual inaccuracies by integrating a retrieval mechanism that grounds responses in relevant information [lewis2020rag].
I formally define the problem as follows. Let $q$ represent the query, $c$ the context information retrieved from external sources, and $r$ the output sequence generated by the model. The architecture consists of two components: (i) a retriever $p_{\eta}(c|q)$ with parameters $\eta$, which selects a top-$K$ set of documents $k$ based on the query $q$; and (ii) a generator $p_{\theta}(r_{i}|q,c,r_{1:i-1})$, with parameters $\theta$, which generates each token $r_{i}$ conditioned on the query $q$, the retrieved context $c$, and previously generated tokens $r_{1:i-1}$.
@@
This implementation follows the RAG-Sequence variant proposed by lewis2020rag, in which a single retrieved context is used to generate the entire response.
Unlike the RAG-Token approach, which may retrieve a different document for each generated token, the former maintains global coherence and is computationally more efficient. Given the goal of ensuring factual grounding and semantic consistency in healthcare-related question answering, this variant is particularly suitable for the present work. In this study, the retrieval and generation components are implemented separately and combined during inference; the complete workflow is described in Section 4.3.
@@ 3.2 Evaluation Metrics
This section describes the metrics employed to evaluate the quality of the responses generated by the system, referred to as the System Summary (SS), against the ground-truth responses, referred to as the Reference Summary (RS), which in this study correspond to human-authored answers.
The evaluation adopts multiple complementary metrics, including ROUGE, BLEU, METEOR, BERTScore, and NMISS, which jointly assess lexical, semantic, and contextual alignment between SS and RS. For all metrics, both SS and RS are represented as sets of $n$-grams. Formally, let $S_{\text{SS}}=\{s_{1},s_{2},\dots,s_{l}\}$ denote the set of $n$-grams in the SS, and $S_{\text{RS}}=\{r_{1},r_{2},\dots,r_{m}\}$ denote the set of $n$-grams in the RS.
ROUGE [lin2004rouge] evaluates textual overlap in terms of precision, recall, and F1-score, after standard text preprocessing [sarkar2019text]. More formally, ROUGE-$N$ measures the $n$-gram overlap between the SS and RS. The corresponding precision, recall, and F1-score are defined as:
@@
$\text{F1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (4)
where $|S_{\text{SS}}\cap S_{\text{RS}}|$ denotes the number of overlapping $n$-grams, $|S_{\text{SS}}|$ the total number of $n$-grams in the SS, and $|S_{\text{RS}}|$ the total number of $n$-grams in the RS. Precision measures the factual accuracy of the generated text, while recall quantifies the coverage of the reference information. Accordingly, low precision suggests the presence of hallucinations, whereas low recall indicates the presence of missing or incomplete content.
A widely used variant is ROUGE-L, which measures the Longest Common Subsequence (LCS) between the SS and RS. The LCS represents the longest ordered sequence of words appearing in both summaries, without requiring consecutive word matching [cormen2022introduction]. ROUGE-L thus captures sentence-level fluency and structural similarity, complementing the $n$-gram–based metrics.
BLEU [papineni2002bleu] is designed to evaluate the quality of SS responses by comparing them against one or more RS answers. BLEU accounts for variability in word choice and word order found in different, yet valid, translations of the same source sentence. It quantifies the overlap by measuring the number of $n$-grams from the SS that are present in the RS. The core of the BLEU metric is the modified $n$-gram precision, which addresses the over-generation problem by treating each reference word as exhausted once a matching candidate word is found.
@@
The final BLEU score is computed by taking the geometric mean of the modified $n$-gram precisions, up to length $N$ and typically set to 4, and multiplying by the brevity penalty:
@@
METEOR [banerjee2005meteor] addresses several limitations of BLEU by emphasizing word-to-word alignment between the SS and RS, accounting for exact matches, stems, and synonyms. For each SS–RS pair, METEOR finds an alignment of unigrams that maximizes the number of matches subject to one-to-one mapping and minimal crossing. Precision ($P$) and recall ($R$) are computed as the fractions of matched unigrams in the SS and RS, respectively, and combined through a weighted harmonic mean:
$F_{\text{mean}}=\frac{P\cdot R}{\alpha P+(1-\alpha)R},$
where $\alpha$ is typically set to 0.9.
To capture word order and fragmentation, METEOR groups matched unigrams into the fewest possible number of contiguous chunks such that the words in each chunk are adjacent and in the same order in both SS and RS. Let $T$ denote the number of such chunks, and $|V|$ the total number of matched unigrams. The fragmentation penalty is then computed as:
$\text{Penalty}=0.5\left(\frac{T}{|V|}\right)^{3}.$
The final METEOR score is given by:
$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{Penalty}).$
This formulation penalizes disordered or fragmented alignments, ensuring that longer, well-ordered matches contribute more to the final score.
BERTScore is an evaluation metric for text generation that leverages pre-trained BERT embeddings [zhang2019bertscore]. Unlike traditional metrics that rely on exact $n$-gram matches, BERTScore measures the semantic similarity between tokens in the SS and the RS by comparing their contextual embeddings. This method captures the semantic meaning even when the exact words differ. It generates contextual embeddings for tokens in both the SS and RS, where each embedding reflects the surrounding context of the word. The similarity between tokens is computed using cosine similarity. For each token in the RS, BERTScore identifies the most similar token in the SS, and vice versa, to compute precision and recall.
Formally, given a RS sentence $x=\{x_{1},\ldots,x_{k}\}$ and a SS sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$, the token embeddings are represented as $\{e(x_{1}),\ldots,e(x_{k})\}$ and $\{e(\hat{x}_{1}),\ldots,e(\hat{x}_{l})\}$, where $e(\cdot)$ represents the contextual embedding for each token. The similarity between two embeddings is computed using cosine similarity. Precision ($P_{\text{BERT}}$) and recall ($R_{\text{BERT}}$) are calculated as follows:
@@
3.3 Negative Missing Information System Scoring
@@ 3.3.1 Rationale
The metrics discussed above evaluate the quality of machine-generated responses by measuring their lexical or semantic overlap with reference answers. While effective, these approaches overlook the fact that non-matching words or $n$-grams may still carry meaningful information if they align with the broader context of the question or source material.
In practice, LLMs often extend their responses beyond the explicit instructions, adding supplementary content, as illustrated in Table 1. Such additions can take two forms: i) hallucinations, where the model introduces factually incorrect or irrelevant information, and ii) contextually relevant elaborations, where the model enriches the answer by incorporating information consistent with the provided context, as in RAG-based generation.
Recognizing this distinction is essential for a more accurate evaluation of model behavior. By explicitly accounting for non-matching elements, it becomes possible to distinguish between genuinely hallucinated content and valuable, contextually appropriate extensions. The NMISS captures the potential informational value of these unmatched words by assessing their alignment with the contextual evidence provided. This rationale acknowledges that additional content in LLM outputs is not inherently erroneous but can instead reflect a form of reasoning or contextual expansion that traditional metrics fail to reward.
@@ 3.3.2 Formulation
Let $\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ represent the set of reference answers, $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$ the set of system-generated answers, and $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ the set of context texts associated with each reference answer. Define the set of matching tokens between system-generated answers and reference answers as $\mathcal{P}(s_{i},r_{i})$, and let $\mathcal{N}(s_{i},r_{i})=s_{i}\setminus\mathcal{P}(s_{i},r_{i})$ represent the tokens in $s_{i}$ that do not match $r_{i}$. Finally, let $\mathcal{A}(s_{i},c_{i})=\mathcal{N}(s_{i},r_{i})\cap c_{i}$, which represents the set of unmatched tokens in $s_{i}$ that match the context $c_{i}$. The function is defined as:
@@
which evaluates how well system-generated answers align with both reference answers and contextual information.
The reference-based evaluation function is defined as follows:
$f_{\text{ref},i}(s_{i},r_{i})=\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$ (13)
while the context-based evaluation function, which computes the score based on tokens matching the context, is defined as:
$f_{\text{cxt},i}(s_{i},c_{i})=\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$ (14)
The NMISS evaluation function is given by:
@@
To preserve the integrity of standard scoring metrics, NMISS introduces two key mechanisms: a weighted scheme and a maximum function. These mechanisms jointly balance reference-based and context-based contributions, mitigating potential distortions such as score inflation or deflation when contextual information is incorporated.
The weighted scheme harmonizes the relative influence of reference overlap and contextual alignment. Because contextual passages are typically much longer and information-rich than reference responses, a simple averaging would overemphasize the contextual component. To address this, the weights $\lambda_{1}$ and $\lambda_{2}$ regulate their respective contributions, ensuring that neither dominates the final score. This mechanism enables NMISS to recognize cases where a system-generated response extends beyond the reference while remaining coherent and faithful to the retrieved evidence.
The maximum function serves as a safeguard against disproportionate score drops caused by denominator inflation in recall-sensitive formulations. This situation often arises with verbose models that rely on contextual paraphrasing when uncertain. Although such verbosity may not introduce hallucinations, it can generate noise—information that adds limited value relative to the reference. By applying a floor, NMISS reverts to the reference-based score whenever the context-based contribution disproportionately lowers the metric due to a large $\lambda_{2}$ term in the denominator. This ensures rewarding genuine contextual relevance while preventing technical artifacts from suppressing the overall score.
This scoring framework is particularly advantageous in question-answering tasks, where model responses may diverge lexically from the reference yet remain contextually valid. The detailed implementation is provided in Algorithm 1.
@@
[…]
@@
8:  $\mathcal{P}(s_{i},r_{i})\leftarrow s_{i}\cap r_{i}$ $\triangleright$ Matched tokens
9:  $\mathcal{N}(s_{i},r_{i})\leftarrow s_{i}\setminus\mathcal{P}(s_{i},r_{i})$ $\triangleright$ Unmatched tokens
@@
13:  $\mathcal{A}(s_{i},c_{i})\leftarrow\mathcal{N}(s_{i},r_{i})\cap c_{i}$ $\triangleright$ Context-matched tokens
@@
To formally evaluate NMISS outperformance, let $\mathcal{H}=\{h_{1},h_{2},\dots,h_{n}\}$ denote the set of hallucination flags associated with the model-generated responses $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{n}\}$. Each flag $h_{i}\in\{0,1\}$ indicates whether response $s_{i}$ is hallucinated ($h_{i}=1$) or not ($h_{i}=0$). NMISS is computed only over non-hallucinated responses. Formally, let
$I=\{1,2,\dots,n\},\quad I_{\text{NH}}=\{\,i\in I\;\mid\;h_{i}=0\,\}$
From this subset, I extract the valid cases where the traditional metric score $f_{\text{ref},i}(s_{i},r_{i})$ lies strictly between two thresholds $\tau_{\min}$ and $\tau_{\max}$, with $0\leq\tau_{\min}<\tau_{\max}\leq 1$. This constraint filters out trivial cases, which are perfect matches or complete mismatches:
$I_{\text{valid}}=\{\,i\in I_{\text{NH}}\;\mid\;\tau_{\min}<f_{\text{ref},i}(s_{i},r_{i})<\tau_{\max}\,\}$
The corresponding subset of responses is then defined as
$\mathcal{S}_{\text{valid}}=\{\,s_{i}\;\mid\;i\in I_{\text{valid}}\,\}$
where $f_{\text{ref},i}(s_{i},r_{i})$ denotes the score computed by a traditional reference-based metric (see Stage 1 in Algorithm 1).
For each valid case $i\in I_{\text{valid}}$, NMISS is considered to outperform the traditional metric if
$f_{\text{NMISS},i}(s_{i},r_{i},c_{i})>f_{\text{ref},i}(s_{i},r_{i})$
where $f_{\text{NMISS},i}$ is the NMISS score calculated over the triplet $(s_{i},r_{i},c_{i})$, incorporating both reference alignment and contextual grounding.
The overall percentage of NMISS outperformance across all valid cases is thus computed as
$\frac{\big|\{\,i\in I_{\text{valid}}\;\mid\;f_{\text{NMISS},i}(s_{i},r_{i},c_{i})>f_{\text{ref},i}(s_{i},r_{i})\,\}\big|}{|I_{\text{valid}}|}\times 100$ (16)
This formulation maintains full consistency with the operational definition of NMISS and isolates its contribution precisely where traditional metrics underperform. By focusing on ambiguous, semantically rich cases that lie outside trivial bounds, NMISS outperformance becomes a robust indicator of a model’s ability to generate contextually grounded yet non-literal responses.
@@ 4.1 Data
Data comprises text-based news in Italian, collected from various sources, and all focused on the health domain within Italy. This collection encompasses regional updates, significant scientific developments, national headlines, and news on research and development, providing comprehensive coverage of health-related issues nationwide.
News dataset consists of 126,470 articles from 2010 to 2024; their frequency is shown in Figure 1. A pre-processing step was employed to clean the raw text of HTML tags and other unwanted artifacts to ensure data quality. Token statistics are calculated to understand the overall structure of the data and are shown in Table 2. On average, each article contains approximately 348 tokens, with a standard deviation of 264 tokens, indicating some variability in article length. The shortest article consists of just 6 tokens, while the longest contains up to 5,594 tokens. The median article length is 284 tokens, with the interquartile range spanning from 195 to 421 tokens, reflecting a diverse range of article sizes.
@@
4.2 Human Evaluation Protocol
To build a robust evaluation framework, I constructed a dataset of 100 public health questions [^1] The full list of questions is available upon request.. The questions span a range of complexity, from broad, general topics to highly specific or ambiguous ones, yielding varying levels of interpretability. To systematically assess this variation, questions were grouped into three difficulty categories—low, mid, and high—based on average character length, used as a proxy for cognitive and semantic complexity [^2] Respectively, low: 30 characters; mid: 40 characters; high: 50 characters.. Each question was reviewed by a domain expert to ensure relevance and factual accuracy, and a manual reference answer is prepared for every entry. Table 3 presents sample questions by category.